Could someone please help?
I am suppose to solve proportion
28/16=x-8/20

Answer by stanbon(75887)

solve proportion
28/16 = (x-8)/20
----
(7/4) = (x-8)/20
---
Multiply both sides by 20 to get:
35 = x-8
x = 43
================
Cheers,
Stan H.